OpenStudy (anonymous):

Solve d^2y/dx^2 = a/(b-y)^2

OpenStudy (unklerhaukus):

separate the variables

OpenStudy (unklerhaukus):

\[\frac{\text d^2y}{\text dx^2} =\frac a{(b-y)^2}\] \[{(b-y)^2}\cdot{\text d^2y} = a\cdot\text dx^2\] \[\iint{(b-y)^2}\cdot{\text d^2y} = \iint a\cdot\text dx^2\]

OpenStudy (anonymous):

it is \[d^2y \] not \[dy^2\] right??

OpenStudy (unklerhaukus):

right ,

OpenStudy (anonymous):

then how can we double integrate?